# Premise-Augmented Reasoning Chains Improve Error Identification in Math Reasoning with LLMs

### v4

@@ 1 Introduction
Chain-of-thought reasoning enhances problem solving by breaking down complex tasks into a series of logical steps, improving the accuracy and clarity of decision-making processes. There has been a well-documented success of the reasoning capabilities of large language models (LLMs) with chain of thought (CoT; Wei et al. (2023)) across applications such as embodied reasoning (Gao et al., 2023; Philipov et al., 2024), code generation (Jiang et al., 2024), and mathematical and scientific reasoning (Imani et al., 2023; Ahn et al., 2024). However, the corresponding evaluation methods and metrics have a significant limitation: they focus solely on the correctness of the final answer, neglecting intermediate reasoning processes and rationales that contribute to it. Although the final answer serves as a proxy for the reasoning capability, it is not sufficient to judge the reasoning performance of the model. Hence, evaluating the final answer provides a narrower view of the reasoning capabilities. Since the intermediate reasoning process is equally important as getting the correct final answer (Huang & Chang, 2023; Golovneva et al., 2023; Prasad et al., 2023), a comprehensive evaluation of the reasoning chains is crucial to holistically understanding the reasoning capabilities of LLMs.
Evaluation of a reasoning chain has been studied in literature in the context of self-verification (Weng et al., 2023). Previous work has shown that LLMs struggle to find reasoning errors in chain-of-thought traces without the help of external verifiers (Stechly et al., 2024; Wu et al., 2024; Tyen et al., 2024), casting doubt on the overoptimism of LLMs’ self-critique abilities. Existing research focusing on the evaluation of reasoning chains in LLMs can be broadly categorized into reference-based and reference-free methods. Reference-based methods, which rely on the availability of a ground truth reasoning chain (Welleck et al., 2021; Han et al., 2024; Tian et al., 2021), are reliable but are constrained by the significant cost of human annotations, restricting their application. In contrast, reference-free methods (Prasad et al., 2023; Golovneva et al., 2023; Zhu et al., 2024) bypass the need for annotations, but suffer from two major drawbacks: (1) most such works assign only a chain-level score rather than localizing specific errors, and (2) they often need fine-tuning of task-specific models, restricting their generalizability. One workaround could be using formal proof assistants like Lean (Moura & Ullrich, 2021), which natively support the verifiability of generated proofs (Yang et al., 2024, 2023; Murphy et al., 2024), but this requires a challenging auto-formalization of natural language text (Wu et al., 2022) and often assumes the solution is already known for proof construction.
@@
We restructure conventional linear reasoning chains (LRCs) into PARC (Premise-Augmented Reasoning Chains), by identifying the premises of each step. Identifying those premises improves the traceability of a reasoning chain. We create a corresponding dataset, called PERL (Premises and ERrors identification in Language models) to test LLMs’ capability in identifying premises as well as the effectiveness of premise augmentation in annotating both premises and errors. Additionally, we refine the error taxonomy (as proposed in (Golovneva et al., 2023)) for mathematical reasoning by introducing a new category: “accumulation error.” This error arises when a reasoning step is correct in isolation but is derived from flawed premises (refer to step 5 in Fig.1), resulting in error propagation throughout the chain; while such errors are common, existing work (Lightman et al., 2023; Zheng et al., 2024; Daheim et al., 2024) has not addressed them beyond discarding all subsequent steps once the first mistake appears. Distinguishing accumulation errors from inherently flawed steps (mathematical errors or logical inconsistencies) is crucial for a holistic evaluation of CoT traces. Moreover, establishing explicit premise links between steps allows each step to be verified only under its premises which reduces irrelevant information for each verification step, thus improving the error detection for mathematical problem solving.
@@
The rest of the paper is structured as follows: in Section 2, we discuss relevant prior work in context of mathematical reasoning and verification of reasoning chains with LLMs. Section 3 introduces the research questions, alongside the necessary background as well as mathematical definition of premises, conversion of LRC to PARC, as well as how we identify errors with premises. Section 4 details PERL, our dataset to show effectiveness of the framework, and lastly, in section 5 we discuss the results and insights from our experiments.
@@ 3.1 Premise Augmented Reasoning Chains (PARC)
where the reasoning chain $\mathbf{R}$ can be decomposed into intermediate steps as:
@@ 3.3.2 Error Identification
Our proposed approach refines the classification process by situating each reasoning step $\mathbf{s}_{k}$ explicitly within the context of its premises $\mathcal{P}_{k}$. Algorithm 1 illustrates the overall pipeline of conversion of LRC to PARC and step-by-step error classification. The prompts for the following are reported in Appendix 14 and 15.
@@ 4 PERL: Premises and ERrors identification in Language models
In order to generate the reasoning chains, we used two popular benchmarks (i) GSM8K (Cobbe et al., 2021), a collection of 8,500 grade school math word problems, and (ii) MATH (Hendrycks et al., 2021b), a dataset of 12,500 challenging competition level math word problems. In addition, we use the (iii) Orca-Math dataset by Mitra et al. (2024), a synthetic dataset of 200K math problems alongside solutions written by GPT4Turbo, and the (iv) MetaMathQA dataset by Yu et al. (2023). We first randomly sampled 1000 examples from the GSM8k and MATH test split and the Orca-Math and MetaMathQA training split (since these are training datasets). We began by using the open weight Large Language Model Llama-3.1-8B-Instruct (Grattafiori et al., 2024) to generate step-by-step reasoning chains. Next, we categorized the reasoning chains as positive or negative, depending on whether the final answer matches the ground-truth answer. Next, we randomly sampled 50 positive (correct) and 50 negative (incorrect) reasoning chains. To expand our dataset, we employed GPT-4o to systematically introduce mathematical or logical errors into the correct reasoning chains, creating additional synthetic negative examples (since prior work also focused on such synthetic negatives). When introducing errors, we ensured that the subsequent steps were appropriately modified to reflect the impact of the initial error. In contrast with existing works such as Golovneva et al. (2023), where the perturbation in a step is not followed up in subsequent steps, our synthetic negatives are more realistic.
@@ 5.1 Premise Mapping
Table 3: Accuracy for Error identification of various models under Full Context (baseline) and Premises settings across different datasets. Note that Neg, Syn and Pos means Negative, Synthetic Negatives and Positive splits of our PERL, as explained earlier. For each dataset, we highlighted the models that benefit the most from the PARC.
@@
We note that the scale of the model is an important factor. The increase in recall varies from 8.2% to 32.2% when we scale up for the same model class, and the pattern is consistent. Tables 7, 8, 10 and 9 in the appendix contain the detailed metrics across the splits Positive, Negative and Synthetic negatives for GSM8K, MATH, OrcaMath and MetaMathQA respectively.
@@ A.2 Data statistics
This results in a total of 607 reasoning chains with 203 positives, 214 negatives, and 190 synthetic negatives. In total, we have 2,134 steps annotated as Correct, 321 steps as Mathematical Error, 443 steps as Logical Inconsistency, and 741 steps as Accumulation Error, indicating that native and accumulation errors appear at almost equal rates. Additionally, we model each chain-of-thought as a directed acyclic graph (PARC) based on the step-level premise annotations. Across 1,370 PARCs, each chain contains on average 7.30 steps, representing the length of the chain-of-thought; 11.27 premises, corresponding to the total premise references across steps; and 10.42 edges linking premises to conclusions. The average depth of the PARC is 6.02, measuring the longest path of dependencies, while the maximum width is 1.90, quantifying how many steps can appear at the same layer in the PARC. Finally, the branching factor is 1.37, which is the ratio of edges to nodes, indicating that each step typically spawns fewer than two subsequent steps on average.
@@ A.5 Prompt for Error identification
The prompts used for the baseline approach are shared in Appendix 12 and 13 . The evaluation for our error identification with premises are done with the prompts outlined in Tables 14 and 15,

### v6

@@ 1 Introduction
Chain-of-thought reasoning enhances problem solving by breaking down complex tasks into a series of logical steps, improving the accuracy and clarity of decision-making processes. There has been a well-documented success of the reasoning capabilities of LLMs with chain of thought reasoning (CoT; (Wei et al., 2023)) across applications such as embodied reasoning (Gao et al., 2023; Philipov et al., 2024), code generation (Jiang et al., 2024), and mathematical and scientific reasoning (Imani et al., 2023; Ahn et al., 2024). However, the corresponding evaluation methods and metrics have a significant limitation: they focus solely on the correctness of the final answer, neglecting intermediate reasoning processes and rationales that contribute to it. Although the final answer serves as a proxy for the reasoning capability, it is not sufficient to judge the reasoning performance of the model. Hence, evaluating the final answer provides a narrower view of the reasoning capabilities. Since the intermediate reasoning process is equally important as getting the correct final answer (Huang & Chang, 2023; Golovneva et al., 2023; Prasad et al., 2023), a comprehensive evaluation of the reasoning chains is crucial to holistically understanding the reasoning capabilities of LLMs.
Evaluation of reasoning chains has been studied in literature in the context of self-verification (Weng et al., 2023). Previous work has shown that LLMs struggle to find reasoning errors in chain-of-thought traces without the help of external verifiers (Stechly et al., 2024; Wu et al., 2024; Tyen et al., 2024), casting doubt on the overoptimism of LLMs’ self-critique abilities. Existing research focusing on the evaluation of reasoning chains in LLMs can be broadly categorized into reference-based and reference-free methods. Reference-based methods, which rely on the availability of a ground truth reasoning chain (Welleck et al., 2021; Han et al., 2024; Tian et al., 2021), are reliable but are constrained by the significant cost of human annotations, restricting their application. In contrast, reference-free methods (Prasad et al., 2023; Golovneva et al., 2023; Zhu et al., 2024) bypass the need for annotations, but suffer from two major drawbacks: (1) most such works assign only a chain-level score rather than localizing specific errors, and (2) they often need fine-tuning of task-specific models, restricting their generalizability. One workaround could be using formal proof assistants like Lean (Moura & Ullrich, 2021), which natively support the verifiability of generated proofs (Yang et al., 2024, 2023; Murphy et al., 2024), but this requires a challenging auto-formalization of natural language text (Wu et al., 2022) and often assumes the solution is already known for proof construction.
@@
We restructure conventional linear reasoning chains (LRCs) into PARC (Premise-Augmented Reasoning Chains), by identifying the premises of each step. Identifying those premises improves the traceability of a reasoning chain. We create a corresponding dataset, called PERL (Premises and ERrors identification in Language models) to test LLMs’ capability in identifying premises as well as the effectiveness of premise augmentation in annotating both premises and errors. Additionally, we refine the error taxonomy (as proposed in Golovneva et al. (2023)) for mathematical reasoning by introducing a new category: “accumulation error.” This error arises when a reasoning step is correct in isolation but is derived from flawed premises (refer to step 5 in Fig.1), resulting in error propagation throughout the chain; while such errors are common, existing work (Lightman et al., 2023; Zheng et al., 2024; Daheim et al., 2024) has not addressed them beyond discarding all subsequent steps once the first mistake appears. Distinguishing accumulation errors from inherently flawed steps (mathematical errors or logical inconsistencies) is crucial for a holistic evaluation of CoT traces. Moreover, establishing explicit premise links between steps allows each step to be verified only under its premises which reduces irrelevant information for each verification step, thus improving the error detection for mathematical problem solving.
@@
The rest of the paper is structured as follows: in Section 2, we discuss relevant prior work in context of mathematical reasoning and verification of reasoning chains with LLMs. Section 3 introduces the research questions, alongside the necessary background as well as mathematical definition of premises, conversion of LRC to PARC, as well as how we identify errors with premises. Section 4 details PERL, our dataset to show effectiveness of the framework, and lastly, in section 5 we discuss the results and insights from our experiments.  [^1] Our code and data is available on https://github.com/SagnikMukherjee/PARC
@@ 3.1 Premise Augmented Reasoning Chains (PARC)
where the reasoning chain $\mathbf{r_{\leq t}}$ can be decomposed into intermediate steps as:
@@ 3.3.2 Error Identification
Our proposed approach refines the classification process by situating each reasoning step $\mathbf{s}_{k}$ explicitly within the context of its premises $\mathcal{P}_{k}$. Algorithm 1 illustrates the overall pipeline of conversion of LRC to PARC and step-by-step error classification. The prompts for the following are reported in Appendix 16 and 17.
@@ 4 PERL: Premises and ERrors identification in Language models
In order to generate the reasoning chains, we used two popular benchmarks (i) GSM8K (Cobbe et al., 2021), a collection of 8,500 grade school math word problems, and (ii) MATH (Hendrycks et al., 2021b), a dataset of 12,500 challenging competition level math word problems. In addition, we use the (iii) Orca-Math dataset by Mitra et al. (2024), a synthetic dataset of 200K math problems alongside solutions written by GPT4Turbo, and the (iv) MetaMathQA dataset by (Yu et al., 2023). We first randomly sampled 1000 examples from the GSM8k and MATH test split and the Orca-Math and MetaMathQA training split (since these are training datasets). We began by using the open weight Large Language Model Llama-3.1-8B-Instruct (Grattafiori et al., 2024) to generate step-by-step reasoning chains. Next, we categorized the reasoning chains as positive or negative, depending on whether the final answer matches the ground-truth answer. Next, we randomly sampled 50 positive (correct) and 50 negative (incorrect) reasoning chains. To expand our dataset, we employed GPT-4o to systematically introduce mathematical or logical errors into the correct reasoning chains, creating additional synthetic negative examples (since prior work also focused on such synthetic negatives). When introducing errors, we ensured that the subsequent steps were appropriately modified to reflect the impact of the initial error. In contrast with existing works such as Golovneva et al. (2023), where the perturbation in a step is not followed up in subsequent steps, our synthetic negatives are more realistic.
@@ 5.1 Premise Mapping
Table 3: Accuracy for Error identification of various models under Full Context (baseline) and Premises settings across different datasets. Note that Neg, Syn and Pos means Negative, Synthetic Negatives and Positive splits of PERL, as explained earlier. For each dataset, we highlighted the models that benefit the most from the PARC.
@@
We note that the scale of the model is an important factor. The increase in recall varies from 8.2% to 32.2% when we scale up for the same model class, and the pattern is consistent. Tables 9, 10, 12 and 11 in the appendix contain the detailed metrics across the splits Positive, Negative and Synthetic negatives for GSM8K, MATH, OrcaMath and MetaMathQA respectively.
@@ Results:
5.3 Experiments on ProcessBench
Beyond the PERL dataset, we also evaluate our framework on ProcessBench (Zheng et al., 2024), a benchmark that contains human-annotated stepwise labels for mathematical solutions. Unlike PERL, ProcessBench only includes annotations up to the first incorrect step in each solution.
Experimental Setup.
We focus on the step-level error identification task, following the protocol used by Zheng et al. (2024). Specifically, for correct solutions (i.e., ones where the final answer is correct), every step is marked as correct. For incorrect solutions, the task is to identify the first incorrect step in the chain; all subsequent steps are automatically deemed irrelevant by design of the benchmark. We adopt the same prompting strategy as in our previous experiments, with LLMs classifying each step as correct or incorrect. The baseline is a standard critic model as proposed by (Zheng et al., 2024) that identifies the first erroneous step from the solution as a whole.
Results and Observations.
Tables 6 and 7 show that our findings on ProcessBench (on the GSM8K and MATH splits) align with the trends we observe on PERL. Premise-augmented verification continues to outperform verification conducted in the full reasoning chain. We see consistent gains from evaluating each step under its extracted premises and classifying its correctness accordingly. This further strenghtens our previous findings.
Model Method Correct Wrong F1 Delta
Qwen 2.5 7B Baseline 66.3 36.7 47.2
Ours 60.1 38.6 47.0 -0.2
Qwen 2.5 32B Baseline 97.9 43.0 59.8
Ours 95.9 55.1 70.0 +10.2
Qwen 2.5 72B Baseline 98.4 61.4 75.6
Ours 97.8 59.7 74.1 -1.5
Llama 3.1 8B Baseline 17.1 36.7 23.3
Ours 33.7 37.8 35.6 +12.3
Llama 3.1 70B Baseline 77.7 57.5 66.1
Ours 89.6 70.0 78.6 +12.5
Table 6: Performance comparison on the GSM8K split of ProcessBench with Baseline vs. PARC (Ours implies premise augmented verification) for various model.
Model Method Correct Wrong F1 Delta
Qwen 2.5 7B Baseline 46.0 25.4 32.7
Ours 45.6 41.2 43.3 +10.6
Qwen 2.5 32B Baseline 90.0 22.4 35.9
Ours 86.9 53.9 66.5 +30.7
Qwen 2.5 72B Baseline 88.5 33.7 48.8
Ours 86.7 53.9 66.5 +17.7
Llama 3.1 8B Baseline 5.6 19.1 8.7
Ours 11.0 27.5 15.7 +7.1
Llama 3.1 70B Baseline 32.4 32.8 32.6
Ours 61.6 55.4 58.3 +25.7
Table 7: Performance comparison on the MATH split of ProcessBench with Baseline vs. PARC (Ours implies premise augmented verification) for various model.
@@ A.2 Data statistics
This results in a total of 607 reasoning chains with 203 positives, 214 negatives, and 190 synthetic negatives. In total, we have 2,134 steps annotated as Correct, 321 steps as Mathematical Error, 443 steps as Logical Inconsistency, and 741 steps as Accumulation Error, indicating that native and accumulation errors appear at almost equal rates. Additionally, we model each chain-of-thought as a directed acyclic graph (PARC) based on the step-level premise annotations. Across 607 PARCs, each chain contains on average 7.30 steps, representing the length of the chain-of-thought; 11.27 premises, corresponding to the total premise references across steps; and 10.42 edges linking premises to conclusions. The average depth of the PARC is 6.02, measuring the longest path of dependencies, while the maximum width is 1.90, quantifying how many steps can appear at the same layer in the PARC. Finally, the branching factor is 1.37, which is the ratio of edges to nodes, indicating that each step typically spawns fewer than two subsequent steps on average.
@@ A.5 Prompt for Error identification
The prompts used for the baseline approach are shared in Appendix 14 and 15 . The evaluation for our error identification with premises are done with the prompts outlined in Tables 16 and 17,